OpenStudy (anonymous):

Solve the system of equations. 6x + 2y = 10 and -6x - 2y = 12 A. (0, 5) B. (-1, 8) C. (2, -1) D. no solution E. infinitely many solutions

hero (hero):

no solution

OpenStudy (hoblos):

D

OpenStudy (anonymous):

no solution. Because when you add the 2 equations together you get 0=2

OpenStudy (anonymous):

2y= -6x+10 2y/2= -6x/2 +10/2 y= -3x + 5

OpenStudy (anonymous):

sorry that should be 0=22

OpenStudy (anonymous):

-6x-2y = 12 -2y = 6x+ 12 -2y/-2 = 6x/-2 +12/-2 y= -3x + -6